An account pays 8.4% compounded continuously. What is the effective quarterly rate? Enter your answer as a percentage without the % sign (e.g., enter 2.05% as 2.05). (Use formulas and Step-by-Step NOT excel)
Answer #1

Effective annual rate with continuous compounding  = ert = e8.4% -1 = 8.76%

Since effective rate with quarterly compounding and effective rate with continuous compounding rates are same

(1+Effective quarterly Rate)4-1 = 8.76%
Effective Quarterly Rate = (1+8.76%)1/4-1 = 2.12%

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well